A researcher orders a solution of 32.5% glucose for her lab. However, she needs a stronger solution, one that is 46% glucose. Fortunately, she has 15.6 liters of 91.9% glucose solution in the stock room. Assuming there is an adequate supply of the 32.5% solution, how many liters of 46% solution can be made by mixing the two together?
Found 2 solutions by lwsshak3, mananth:
Answer by lwsshak3(11628)

researcher orders a solution of 32.5% glucose for her lab. However, she needs a stronger solution, one that is 46% glucose. Fortunately, she has 15.6 liters of 91.9% glucose solution in the stock room. Assuming there is an adequate supply of the 32.5% solution, how many liters of 46% solution can be made by mixing the two together?
..
let x=liters of 32.5% solution to be used
x+15.6=liters of 46% solution that can be made
..
32.5%(x)+91.9%(15.6)=46%(x+15.6)
.325x+14.34=.46x+7.18
-.135x=-7.16
x=53.04 liters of 32.5% solution used in mixture
53.04+15.6=68.64
ans:
68.64 liters of 46% solution can be made


Answer by mananth(16946)

percent ---------------- quantity
91.9 ---------------- 15.6
32.5 ---------------- x
46 ---------------- 15.6+x
...
91.9*15.6+32.5*x =46(15.6 +x)
1433.64 +32.5x=717.6+46x
32.5x-46 x=717.6-1433.64
91.9*15.6+32.5*x =46(15.6 +x)
1433.64 +32.5x=717.6+46x
32.5x-46x=717.6-1433.64
-13.5x=-716.04
/-13.5
x=53.04
53.4+15.6=69 liters of 46%